OpenStudy (anonymous):

3 over x-2 + 3 over x +2 = 2

OpenStudy (anonymous):

\[\frac{ 3 }{ x-2 }+\frac{ 3 }{x+2}=2\]

OpenStudy (anonymous):

Develop the equation: 3(x + 2) + 3(x - 2) = 2(x^2 - 4) 6x = 2x^2 - 8 -> x^2 - 3x - 4 = 0. Solve the quadratic equation.